# Structure of activity in multiregion recurrent neural networks

### v1

@@ V.2 Stability
where the factor $1/u^{\mu}$ is inspired by the generalized symmetry property. We will show that $\partial_{t}E\leq 0$ for all $\sigma^{\mu\nu}$ when $S^{\mu\nu}_{0}$ is in a configuration claimed to be stable. Computing the time derivative and subsequently replacing $\sigma^{\mu\nu}\rightarrow u^{\mu}\sigma^{\mu\nu}$ gives $\partial_{t}E=\sum_{\nu}e^{\nu}$, where
$e^{\nu}=$ $-\sum_{\mu}(\sigma^{\mu\nu})^{2}+b^{\nu}\psi_{0}^{\nu}\sum_{\mu}\sigma^{\nu\mu}\sigma^{\mu\nu}$
$+2b^{\nu}\psi^{\prime\nu}_{0}\Big{(}\sum_{\mu}S_{0}^{\nu\mu}\sigma^{\mu\nu}\Big{)}\Big{(}\sum_{\rho}S_{0}^{\nu\rho}\sigma^{\nu\rho}\Big{)}$
$+h^{\nu}(\sigma^{\nu\nu})^{2}\left(\psi_{0}^{\nu}+2\psi^{\prime\nu}_{0}(S_{0}^{\nu\nu})^{2}\right),$ (26)
@@
where $\psi^{\nu}_{0}=\psi(A^{\nu})$ and $\psi^{\prime\nu}_{0}=\psi^{\prime}(A^{\nu})$. We perform a symmetric-antisymmetric decomposition, ${\sigma^{\mu\nu}=\alpha^{\mu\nu}+\beta^{\mu\nu}}$, where ${\alpha^{\mu\nu}=\alpha^{\nu\mu}}$ and ${\beta^{\mu\nu}=-\beta^{\nu\mu}}$, and discard the term $-2\sum_{\mu}\alpha^{\mu\nu}\beta^{\mu\nu}$, which vanishes under the outer $\nu$ sum. We then have for a region $\nu$ in routing mode
$e^{\nu}=2b^{\nu}\psi^{\prime\nu}_{0}\Big{(}\sum_{\mu}S_{0}^{\mu\nu}\alpha^{\mu\nu}\Big{)}^{2}+h^{\nu}\psi_{0}^{\nu}(\alpha^{\nu\nu})^{2}\\ -2\sum_{\mu}(\beta^{\mu\nu})^{2}-2b^{\nu}\psi^{\prime\nu}_{0}\Big{(}\sum_{\mu}S_{0}^{\mu\nu}\beta^{\mu\nu}\Big{)}^{2}.$ (27)
@@
The first term is nonpositive since $\psi^{\prime\nu}_{0}<0$. The second is nonpositive when $h^{\nu}<0$, i.e., $a^{\nu}<b^{\nu}$. The third and fourth terms, involving $\beta^{\mu\nu}$, are net-nonpositive for all $\beta^{\mu\nu}$ if, and only if, ${-{\psi^{\prime}(\Delta)\Delta}/{\psi(\Delta)}\leq 1}$. This quantity varies between 0 to $1/2$ as $\Delta$ varies from zero to infinity, so this holds. Thus, $\partial_{t}E\leq 0$.
@@
$e^{\nu}=-\left(1-\frac{b^{\nu}}{a^{\nu}}\right)\sum_{\mu\neq\nu}(\alpha^{\mu\nu})^{2}+{2a^{\nu}\psi^{\prime\nu}_{0}(S_{0}^{\nu\nu})^{2}}(\alpha^{\nu\nu})^{2}\\ -\left(1+\frac{b^{\nu}}{a^{\nu}}\right)\sum_{\mu}(\beta^{\mu\nu})^{2}.$ (28)
@@
Conversely, if there is a routing mode with ${a^{\mu}>b^{\mu}}$, we obtain $\partial_{t}E>0$ by picking $\alpha^{\mu\mu}$ to be nonzero and all other components of $\alpha^{\mu\nu}$ and $\beta^{\mu\nu}$ to be zero. Similarly, if there is a non-routing mode with $a^{\mu}<b^{\mu}$, we obtain $\partial_{t}E>0$ by picking $\alpha^{\mu\nu}$ to be nonzero and orthogonal to $S^{\mu\nu}_{0}$ when contracted over $\nu$, and everything else zero. These choices of $\sigma^{\mu\nu}$ indicate directions along which perturbations grow away from the fixed point. For a region $\mu$ in routing mode, there are directions in which the local energy neither grows nor shrinks, $\partial_{t}E=0$, obtained by $\alpha^{\mu\nu}$ being nonzero and orthogonal to $S^{\mu\nu}_{0}$ when contracted over $\nu$, and everything else zero. We show in Sec. V.4 that such directions correspond to translation along a continuous attractor manifold.
@@
While $\partial_{t}E<0$ rigorously indicates (marginal) stability, $\partial_{t}E>0$ does not necessarily indicate instability; it might reflect transient dynamics en route to a stable state. Nevertheless, we find through numerical diagonalization of the Jacobian that the perturbations with $\partial_{t}E>0$ given above indeed represent unstable directions. An interesting, as yet unanswered question is whether this system, under the symmetry constraint, possesses a global energy function that ensures convergence to fixed points from any initial condition, similar to regular neural networks with coupling symmetry.
Before we introduce disorder into the model, we highlight how the method for routing neural signals used here differs from traditional approaches. Traditional approaches typically relied on manipulating individual neurons or synapses through neuromodulation, inhibition, or gain modulation. In contrast, our approach bases the transmission of signals between regions on the geometry of neural population activity. Specifically, instead of significantly reducing $\psi_{0}^{\mu}$ to switch a region $\mu$ to a non-routing mode–akin to gain control–our model leverages the interaction between connectivity geometry and nonlinear recurrent dynamics. This interaction aligns neural activity with subspaces that allow or prevent cross-region communication. This is illustrated in Fig. 3(b). Thus, our model offers a new perspective on routing in neural circuits, emphasizing the excitation of subspaces sculpted by the structure of connectivity and network dynamics.
V.3 Effect of disorder
Figure 5: Structure of activity in networks with disorder and symmetric effective interactions among regions. (a) Relationship between $A^{\mu}$ and $g^{\mu}$ for various values of $b^{\mu}$ in the DMFT. (b) Solutions for the two-point function $\Delta^{\mu}(\tau)$ for the parameter values indicated by the markers in (a). (c–e) are the same as (a–c) in Fig. 4, but with disorder, whose levels are shown in (a). All regions have $g^{\mu}>1$, so regions produce chaotic fluctuations unless tamed by current-based activity. In the leftmost plots, chaos is suppressed in all regions, and all regions are in routing mode. In the middle plots, all regions are in routing mode, and chaotic fluctuations exist alongside the structured current-based activity in region 1. In the rightmost plots, region 1 is in disorder-dominated non-routing mode, and regions 2–5 are in routing mode.
In this section, we introduce disorder into the model by allowing nonzero $g^{\mu}$, potentially leading to high-dimensional chaotic fluctuations. Such fluctuations can disrupt signal transmission between regions. In particular, chaos scrambles neural signals, preventing coherent communication. This is comparable to the structure-dominated non-routing mode discussed above, where structured within-region activity impedes signal flow. Both types of non-routing behaviors exemplify the underlying tension between signal generation and transmission.
The symmetric structure of the effective-interaction tensor implies that, even with disorder, the currents converge to fixed points, $S^{\mu\nu}_{0}$. Due to the possibility of coexisting high-dimensional chaotic fluctuations, $\Delta^{\mu}(t,t^{\prime})$ can depend on time. We seek stationary solutions, $\Delta(t,t^{\prime})=\Delta(\tau)$ where $\tau=t-t^{\prime}$, in which case the DMFT equations become
$\psi_{0}^{\nu}$ $=\psi(\Delta^{\nu}(0)),$ (29b)
$\partial_{\tau}^{2}\Delta^{\mu}(\tau)$ $=-\partial_{\Delta^{\mu}}V^{\mu}(\Delta^{\mu};A^{\mu}),$ (29c)
$V^{\mu}(\Delta^{\mu};A^{\mu})$ $=-\frac{(\Delta^{\mu})^{2}}{2}$
$\hskip 11.99998pt+(g^{\mu})^{2}\Phi(\Delta^{\mu},\Delta^{\mu}(0))+A^{\mu}\Delta^{\mu},$ (29d)
$A^{\mu}$ $=\sum_{\nu}(S_{0}^{\mu\nu})^{2},$ (29e)
where we replaced ${(1+\partial_{t})(1+\partial_{t^{\prime}})\rightarrow 1-\partial_{\tau}^{2}}$ in Eq. 13, then integrated its rhs with respect to $\Delta^{\mu}(\tau)$. To obtain $\Phi(\Delta_{12},\Delta_{11})$, we first define $C(\Delta_{12},\Delta_{11})$ by setting $\Delta_{11}=\Delta_{12}$ in Eq. 15b, then integrate with respect to $\Delta_{12}$.
These equations generalize Eq. V.1 to the case where the squared $L^{2}$-norms of the rows of the current matrix differ from the equal-time two-point functions, ${A^{\mu}\neq\Delta^{\mu}(0)}$, due to chaotic fluctuations contributing to the variance of activity in addition to the currents. As in Sompolinsky et al. [33], $\Delta^{\mu}(\tau)$ acts like a Newtonian particle in a Mexican-hat potential, $V^{\mu}(\Delta^{\mu};A^{\mu})$. The values of $\psi^{\mu}_{0}$ and thus $\Delta^{\mu}(0)$ are determined by $a^{\mu}$ and $b^{\mu}$ as in the disorder-free case, but $A^{\mu}$ is as yet undetermined. We exchange the dependence of the potential on $A^{\mu}$ for a dependence on the large-$\tau$ value of the two-point function, ${\Delta^{\mu}(\infty)=\lim_{\tau\rightarrow\infty}\Delta^{\mu}(\tau)}$, which satisfies $\partial_{\Delta^{\mu}}V^{\mu}(\Delta^{\mu}(\infty))=0$ because the Newtonian particle must come to rest at the top of a hill to obtain a valid decaying two-point function. This condition can be expressed as
We use this to express the potential as $V^{\mu}(\Delta^{\mu};\Delta^{\mu}(\infty))$, eliminating the dependence on $A^{\mu}$. Finally, $\Delta^{\mu}(\infty)$ is determined by energy conservation, $V^{\mu}(\Delta^{\mu}(0);\Delta^{\mu}(\infty))=V^{\mu}(\Delta^{\mu}(\infty);\Delta^{\mu}(\infty))$. $A^{\mu}$ can then be found using Eq. 30, shown in Fig. 5(a), and the full form of $\Delta^{\mu}(\tau)$ is given by integrating the Newtonian dynamics, shown in Fig. 5(b). A similar analysis was done by Mastrogiuseppe and Ostojic [25].
Solutions of these equations have the following structure, depicted in Figs. 5(c–e). When $g^{\mu}$ is small, chaotic fluctuations are not present in region $\mu$, so $\Delta^{\mu}(\tau)=\Delta^{\mu}(0)=\Delta^{\mu}(\infty)$. Routing and non-routing modes behave as in the disorder-free case, namely, $\psi^{\mu}=1/a^{\mu}$ or $\psi^{\mu}=1/b^{\mu}$, with the same constraints on the placement of zeros in the current matrix, and with stability determined by whether $a^{\mu}$ or $b^{\mu}$ is larger. This non-chaotic regime exists for $g^{\mu}>1$, reflecting the suppression of chaos by currents emanating from within the region, if in non-routing mode, or from other regions, if in routing mode. Compared to the disorder-free case, $A^{\mu}$ is reduced, implying that disorder hinders currents. As $g^{\mu}$ increases, there is a phase transition to a regime in which chaotic fluctuations coexist with currents, corresponding to $\Delta^{\mu}(\infty)<\Delta^{\mu}(0)$ with $\Delta^{\mu}(\tau)$ decaying. In both of these phases, disorder-induced reductions of $A^{\mu}$ can change the topology of the set of fixed points, as we show in Sec. V.4.
When $g^{\mu}$ becomes sufficiently large, there is a phase transition to a new type of non-routing mode dominated by disorder. In this case, $\Delta^{\mu}(\tau)$ decays from $\Delta^{\mu}(0)>0$ to $\Delta^{\mu}(\infty)=0$, and $A^{\mu}=0$. The values of $\psi_{0}^{\mu}$ and $\Delta^{\mu}(\tau)$ are no longer determined by $a^{\mu}$ and $b^{\mu}$; instead, $\Delta^{\mu}(\tau)$ follows the solution of [33] as if there were no structured connectivity. This phase differs from the structure-dominated non-routing mode, described in the previous section, in that signal transmission from other regions is thwarted by high-dimensional chaotic fluctuations rather than structured self-exciting activity, resulting in $S^{\mu\mu}_{0}=0$. That is, chaotic fluctuations scramble incoming and outgoing signals to the degree that reliable signal transmission is not possible. These disorder-induced phase transitions are decoupled across regions because the low-rank structure of the cross-region connectivity prevents propagation of high-dimensional chaotic fluctuations.
@@ V.3 Effect of disorder
V.4 Dimension and topology of the attractor manifold
@@
Figure 6: Convex geometry of the attractor manifold in multiregion networks. (a) Cartoon illustration of fracture variables. In a cartoon “feasible region” defined by a line segment, $x_{2}$ is never zero, leading to a binary fracturing of the current-space manifold. (b) Left: Summary of the geometry of the solution polytope, determined by random choices of $\bm{b}$. For each choice, a point is plotted in a plane with axes representing the log-number of vertices and the number of fracture variables, $f$. Different colors indicate different numbers of regions in routing mode, $M$. Center: Configurations of the current submatrix at fixed points for choices of $\bm{b}$ resulting in $f$ fracture variables, with increasing values of $f$ (ascending) and $M=7$. Right: “Barcode” visualization of all vertices of the feasible region, in $\bm{x}$ space, for specific choices of $\bm{b}$. Here, $M=7$, and thus there are $n=21$ variables in the linear program. Barcodes for $f\in\{3,4,5\}$ are displayed (descending). (c) Averaged sorted values of $b^{\mu}$, conditioned on yielding specified numbers of fracture variables $f$. (d) t-SNE visualizations of fixed-point manifolds, in current space, for various numbers of fracture variables $f$ with $M=7$.
@@
By imposing symmetry on the effective-interaction tensor, the multiregion system acts as an attractor network, with the currents (but, if disorder is present, not necessarily neuronal activations) converging to fixed points. These equilibrium states remain unchanged over a timescale significantly longer than that of individual neurons, becoming infinite as $N\rightarrow\infty$. In neuroscience, attractor dynamics have been theoretically and experimentally successful in explaining memory mechanisms involving discrete and continuous variables, as well as in the integration of continuous variables. Models in this category include discrete attractors, such as Hopfield networks [45]; and continuous attractors, such as line [23], ring [46, 47, 48], and toroidal attractors [49, 50]. These various structures are appropriate for different cognitive tasks; for instance, discrete attractors are useful for tasks requiring the retention and recall of specific information, whereas continuous attractors are useful for tasks involving the tracking or integration of ongoing stimuli or movements.
@@
Our analysis thus far focused on the structure of fixed points without considering the dimension and topology of the manifold on which they reside. We now explore the structure of this current-space manifold, which is linearly embedded in neuronal space, through a connection to convex geometry. We demonstrate that the architecture of the multiregion network facilitates a blend of discrete and continuous attractors, potentially useful for tasks that necessitate tracking continuous signals in a context-specific way. Furthermore, the dimension and topology of the manifold can be modified by adjusting the effective interactions, rather than by rewiring the network architecture. We begin by reducing the problem of determining the structure of the current-space manifold to a linear program, then use this linear-program equivalence to describe the manifold’s dimension and topology.
@@
For both the disordered and non-disordered cases, the structure of the current-space manifold is shaped by three constraints on the submatrix of $S^{\mu\nu}_{0}$ restricted to the set of regions $\mathcal{S}_{\text{route}}$ in routing mode: 1) zero on-diagonals; 2) equality constraints on the squared $L^{2}$-norms of rows (involving $A^{\mu}$); and 3) the generalized symmetry property, ${S_{0}^{\mu\nu}u^{\nu}=S_{0}^{\nu\mu}u^{\mu}}$. We encode these constraints as a linear program with a vector of variables $\bm{x}$. These variables correspond to the squared upper-triangular elements of the current submatrix restricted to regions in routing mode,
$\bm{x}=\{(S^{\mu\nu}_{0})^{2}\>|\>\mu,\nu\in\mathcal{S}_{\text{route}},\>\mu<\nu\}.$ (31)
Thus, the size of $\bm{x}$ is ${n=M(M-1)/2}$, where $M=|\mathcal{S}_{\text{route}}|$ is the number of regions in routing mode. Crucially, there is a nonnegativity constraint, $x_{k}\geq 0$ for $k\in\{1,\ldots,n\}$, because the variables represent squared quantities.
There are $M$ linear equality constraints that can be expressed as $\bm{C}\bm{x}=\bm{A}$. Here, $\bm{A}$ has components $A^{\mu}$, where $A^{\mu}$ is the required squared $L^{2}$-norm of row $\mu$ of the current submatrix, and $\bm{C}$ is an $M$-by-$n$ constraint matrix. Each element of $\bm{C}$ is set to unity or the ratio $b^{\mu}/b^{\nu}$ if the element corresponds to an upper- or lower-triangular element of the current submatrix, respectively; otherwise the element is set to zero. For concreteness, in the case of $M=5$, symmetric pairs of elements of the current submatrix can be indexed from $1$ through $n=10$ as
$\begin{pmatrix}\bullet&1&2&3&4\\ 1&\bullet&5&6&7\\ 2&5&\bullet&8&9\\ 3&6&8&\bullet&10\\ 4&7&9&10&\bullet\\ \end{pmatrix},$ (32)
which gives the constraint matrix

### v3

@@ A.4 Stability analysis via the local energy function
Computing the time derivative $dE/dt$ and subsequently replacing $\sigma^{\mu\nu}\rightarrow u^{\mu}\sigma^{\mu\nu}$ gives $dE/dt=\sum_{\nu}e^{\nu}$, where
$e^{\nu}=-\sum_{\mu}(\sigma^{\mu\nu})^{2}+b^{\nu}\psi_{0}^{\nu}\sum_{\mu}\sigma^{\nu\mu}\sigma^{\mu\nu}+2b^{\nu}\psi^{\prime\nu}_{0}\Big{(}\sum_{\mu}S_{0}^{\nu\mu}\sigma^{\mu\nu}\Big{)}\Big{(}\sum_{\rho}S_{0}^{\nu\rho}\sigma^{\nu\rho}\Big{)}+h^{\nu}(\sigma^{\nu\nu})^{2}\left(\psi_{0}^{\nu}+2\psi^{\prime\nu}_{0}(S_{0}^{\nu\nu})^{2}\right),$ (26)
@@
and $\psi^{\nu}_{0}=\psi(A^{\nu})$ and $\psi^{\prime\nu}_{0}=\psi^{\prime}(A^{\nu})$. We perform a symmetric-antisymmetric decomposition, ${\sigma^{\mu\nu}=\alpha^{\mu\nu}+\beta^{\mu\nu}}$, where ${\alpha^{\mu\nu}=\alpha^{\nu\mu}}$ and ${\beta^{\mu\nu}=-\beta^{\nu\mu}}$, and discard the term $-2\sum_{\mu}\alpha^{\mu\nu}\beta^{\mu\nu}$, which vanishes under the outer $\nu$ sum.
We then have for a region $\nu$ in routing mode
$e^{\nu}=2b^{\nu}\psi^{\prime\nu}_{0}\Big{(}\sum_{\mu}S_{0}^{\mu\nu}\alpha^{\mu\nu}\Big{)}^{2}+h^{\nu}\psi_{0}^{\nu}(\alpha^{\nu\nu})^{2}-2\sum_{\mu}(\beta^{\mu\nu})^{2}-2b^{\nu}\psi^{\prime\nu}_{0}\Big{(}\sum_{\mu}S_{0}^{\mu\nu}\beta^{\mu\nu}\Big{)}^{2}.$ (27)
@@
The first term is nonpositive since $\psi^{\prime\nu}_{0}<0$. The second is nonpositive when $h^{\nu}<0$, i.e., $a^{\nu}<b^{\nu}$. The third and fourth terms, involving $\beta^{\mu\nu}$, are net-nonpositive for all $\beta^{\mu\nu}$ if, and only if, ${-{\psi^{\prime}(\Delta)\Delta}/{\psi(\Delta)}\leq 1}$. This quantity varies between 0 to $1/2$ as $\Delta$ varies from zero to infinity, so this holds. Thus, $dE/dt\leq 0$.
@@
$e^{\nu}=-\left(1-\frac{b^{\nu}}{a^{\nu}}\right)\sum_{\mu\neq\nu}(\alpha^{\mu\nu})^{2}+{2a^{\nu}\psi^{\prime\nu}_{0}(S_{0}^{\nu\nu})^{2}}(\alpha^{\nu\nu})^{2}-\left(1+\frac{b^{\nu}}{a^{\nu}}\right)\sum_{\mu}(\beta^{\mu\nu})^{2}.$ (28)
@@
Conversely, if there is a routing mode with ${a^{\mu}>b^{\mu}}$, we obtain $dE/dt>0$ by picking $\alpha^{\mu\mu}$ to be nonzero and all other components of $\alpha^{\mu\nu}$ and $\beta^{\mu\nu}$ to be zero. Similarly, if there is a non-routing mode with $a^{\mu}<b^{\mu}$, we obtain $dE/dt>0$ by picking $\alpha^{\mu\nu}$ to be nonzero and orthogonal to $S^{\mu\nu}_{0}$ when contracted over $\nu$, and everything else zero. These choices of $\sigma^{\mu\nu}$ indicate directions along which perturbations grow away from the fixed point. For a region $\mu$ in routing mode, there are directions in which the local energy neither grows nor shrinks, $dE/dt=0$, obtained by $\alpha^{\mu\nu}$ being nonzero and orthogonal to $S^{\mu\nu}_{0}$ when contracted over $\nu$, and everything else zero. We show in the main text that such directions correspond to translation along a continuous attractor manifold.
@@
While $dE/dt<0$ rigorously indicates (marginal) stability, $dE/dt>0$ does not necessarily indicate instability; it might reflect transient dynamics en route to a stable state. Nevertheless, we find through numerical diagonalization of the Jacobian that the perturbations with $dE/dt>0$ given above indeed represent unstable directions. An interesting, as yet unanswered question is whether this system, under the symmetry constraint, possesses a global energy function that ensures convergence to fixed points from any initial condition, similar to regular neural networks with coupling symmetry.
@@
A.5 Dimension and topology of the attractor manifold in multiregion networks with symmetric interactions
@@
Figure 8: Convex geometry of the attractor manifold in multiregion networks. (a) Cartoon illustration of fracture variables. In this cartoon “feasible region" defined by the black line segment, $x_{2}$ is never zero, leading to a binary fracturing of the current-space manifold. (b) Left: Summary of the geometry of the solution polytope, determined by 10000 random choices of $\bm{b}$ (uniform components over [1, 3]). For each random choice, we calculated the number of fracture variables, $f$, and the log-number of vertices. We show the mean and standard deviation of log-number of vertices for each different number of fracture variables. Different colors indicate different numbers of regions in routing mode, $M$. Error bars show 2 standard deviations. Center: Configurations of the current submatrix at fixed points for choices of $\bm{b}$ resulting in $f$ fracture variables, with increasing values of $f$ (ascending) and $M=7$. Right: “Barcode” visualization of all vertices of the feasible region, in $\bm{x}$ space, for specific choices of $\bm{b}$. Each horizontal row corresponds to a different choice of $\bm{b}$, each vertical column corresponds to the $k$-th entry in the vector of variables $\bm{x}$, where $k\in\{1,\ldots,n\}$. The shading indicates the normalized value of $x_{k}$ at the vertex. Here, $M=7$, and thus there are $n=21$ variables in the linear program. Barcodes for $f\in\{3,4,5\}$ are displayed (descending). (c) Averaged sorted values of $b^{\mu}$, conditioned on yielding specified numbers of fracture variables $f$. (d) t-SNE visualizations of fixed-point manifolds, in current space, for various numbers of fracture variables $f$ with $M=7$.
@@
By imposing symmetry on the effective-interaction tensor, the multiregion system acts as an attractor network, with the currents converging to fixed points. These equilibrium states remain unchanged over a timescale significantly longer than that of individual neurons, becoming infinite as $N\rightarrow\infty$. In neuroscience, attractor dynamics have explained memory mechanisms involving discrete and continuous variables, as well as the integration of continuous variables [36, 24, 37, 58, 59, 60, 61]. Discrete attractors are useful for tasks requiring the retention and recall of specific information, whereas continuous attractors are useful for tasks involving the tracking or integration of ongoing stimuli or movements.
@@
Our analysis thus far has characterized fixed points of the currents without considering the structure of the manifold on which they reside. We now explore this structure through a connection to convex geometry. We show that the architecture of the multiregion network facilitates a blend of discrete and continuous attractors, useful for tasks that necessitate tracking continuous signals in a context-specific way. Furthermore, the dimension and topology of the manifold can be modified by adjusting the effective interactions rather than by rewiring the network architecture. The current-space manifold is linearly embedded in neuronal space, and thus the neuron-space manifold inherits the dimension and topology of the current-space manifold.
@@
We begin by reducing the problem of determining the structure of the current-space manifold to a linear program. For both the disordered and non-disordered cases, the manifold is shaped by three constraints on the submatrix of $S^{\mu\nu}_{0}$ restricted to the set of regions $\mathcal{S}_{\text{route}}$ in routing mode: 1) zero on-diagonals; 2) equality constraints on the squared $L^{2}$-norms of rows (involving $A^{\mu}$); and 3) the generalized symmetry property, ${S_{0}^{\mu\nu}u^{\nu}=S_{0}^{\nu\mu}u^{\mu}}$. We encode these constraints using a vector of variables $\bm{x}$ corresponding to the squared upper-triangular elements of the current submatrix restricted to regions in routing mode, i.e., ${\bm{x}=\{(S^{\mu\nu}_{0})^{2}\>|\>\mu,\nu\in\mathcal{S}_{\text{route}},\>\mu<\nu\}}$. Thus, the number of variables is ${n=M(M-1)/2}$ where $M=|\mathcal{S}_{\text{route}}|$. Crucially, there is a nonnegativity constraint $x_{k}\geq 0$ for $k\in\{1,\ldots,n\}$ because $x_{k}$ represents a squared quantity. There are $M$ linear equality constraints that can be expressed as $\bm{C}\bm{x}=\bm{A}$. Here, $\bm{A}$ has components $A^{\mu}$, where $A^{\mu}$ is the required squared $L^{2}$-norm of row $\mu$ of the current submatrix, and $\bm{C}$ is an $M$-by-$n$ constraint matrix. Each element of $\bm{C}$ is set to unity or the ratio $b^{\mu}/b^{\nu}$ if the element corresponds to an upper- or lower-triangular element of the current submatrix, respectively; otherwise the element is set to zero (see “Concrete example of the linear program construction” below).
The solution set of this linear program, called the feasible region, is a convex polytope. Barring fine tuning, its dimension $d$ is the number of variables minus the number of constraints,
$d=n-M=\frac{M(M-3)}{2}.$ (29)
This is also the dimension of the current-space manifold. For $M\geq 4$, the manifold is therefore continuous with dimension $d\geq 2$. For $M=3$, the manifold reduces to a zero-dimensional point set. For $M=2$, or for a sufficiently nonuniform constraint vector $\bm{A}$, the linear program is infeasible, i.e., has no solutions. In this case, the assumptions of the linear-program formulation are violated, and the system converges to an exceptional fixed point that can be described analytically by returning to the fixed-point equations (see “Characterization of the exceptional fixed point” below).
To characterize the topology of the current-space manifold in non-exceptional cases, we observe that, for a given point $\bm{x}$ on the feasible region, there are $2^{n}$ corresponding points in current space. This multiplicity arises from the different ways one can choose the signs of the currents. The connectedness, or lack thereof, of the manifold hinges on whether $x_{k}=0$ for each $k\in\{1,\ldots,n\}$ is included in the feasible region. If included, positive and negative current-space branches connect; otherwise, a binary fracture of the manifold is induced. We refer to variables $x_{k}$ that never take on zero as fracture variables and denote their number by $f$. This is visualized in Fig. 8a. Each fracture variable contributes one binary split to the manifold, resulting in $2^{f}$ connected components. Zeros in components of $\bm{x}$ occur only at vertices of the feasible region, so to identify all fracture variables, it suffices to enumerate all vertices.
To generate realizations of this linear program, we first pick a vector $\bm{b}$ by sampling its components $b^{\mu}$ uniformly over $1\leq b^{\mu}\leq 3$ and sort them in ascending order for visualization. We set $A^{\mu}=\psi^{-1}(1/b^{\mu})$ for each $\mu$, assuming the disorder-free case with all regions in routing mode. Using the double-description method of Motzkin [62, 63], we identify all vertices. We plot $f$ against the log-number of vertices for realizations of the linear program with $M\in\{5,6,7\}$ (Fig. 8b, left), finding that the number of vertices grows exponentially with $M$ (Fig. 8b, left). $f$ is negatively correlated with vertex count and is at most $M-1$. Except for $f=3$, fracture variables correspond to currents for the region with largest $b^{\mu}$ (Fig. 8b, center). For $f=3$, there is an additional configuration involving all currents between the three regions with the largest values of $b^{\mu}$. We visualized all $n=21$ vertices for example realizations with $M=7$ and $f\in\{3,4,5\}$ (Fig. 8b, right). Choices of $\bm{b}$ leading to more fracture variables tend to have nonuniform components (Fig. 8c).
We confirmed that the topology predicted by fracture variables matches that of the current-space manifold. For many samples of $\bm{b}$, we evolved the disorder-free DMFT equations from different initial conditions until convergence to fixed points. We then applied t-SNE nonlinear dimensionality reduction to the collection of fixed points [64]. The number of distinct clusters was $2^{f}$ in all cases, as visualized in Fig. 8d. Each cluster has some spread corresponding to the continuous dimensions of variation on the manifold.
The dimension and topology of the current-space manifold are determined by the number of regions in routing mode and the values of $A^{\mu}$, respectively. These quantities can be changed by adjusting $a^{\mu}$, $b^{\mu}$, and $g^{\mu}$. Doing this adjustment dynamically, e.g., through neuromodulation, provides a way to maintain an attractor manifold with a variable dimension and number of connected components, without the need to construct a completely new network architecture. This adaptability could be advantageous for responding to nonstationary tasks or environmental conditions where the computational demands on the attractor system change rapidly and significantly.
A.6 Concrete example of the linear program construction
To analyze the dimension and topology of attractors in the case of symmetric effective interactions, we solve a linear program of the form $\bm{C}\bm{x}=\bm{A}$. For concreteness, in the case of $M=5$, where $M$ is the number of regions in routing mode, symmetric pairs of elements of the current submatrix can be indexed from $1$ through $n=10$ as
$\begin{pmatrix}\bullet&1&2&3&4\\ 1&\bullet&5&6&7\\ 2&5&\bullet&8&9\\ 3&6&8&\bullet&10\\ 4&7&9&10&\bullet\\ \end{pmatrix}.$ (30)
That is, $x_{1}=(S_{0}^{12})^{2}=(S_{0}^{21})^{2}$, $x_{2}=(S_{0}^{13})^{2}=(S_{0}^{31})^{2}$, and so on. This gives the constraint matrix